You have a laser from a gas station and want to know the wavelength and frequency of the light it produces. You sneak into the physics lab one evening to make use of the double slit screens you know are in the back room.

You select a set of two slits with a separation of 0.12 mm and shine the laser onto the slits. The laser is positioned 52 cm behind the slits. You cast the interference pattern onto the wall of the physics lab a distance of 2.8 meters from the slits. After grabbing a ruler, you find that the distance from the central maximum to the 3rd bright spot over is 4.7 cm.

Determine the following:

The wavelength of the laser in SI units

The frequency of the laser in SI units

The angle (measured from a line running from the center of the slits directly across to the

screen) corresponding to the 2nd dark fringe. (Think carefully about the value for m here)
Answer #1


from the given data

slit separation, d = 0.12 mm = 0.12*10^-3 m

distance from slits to screen, R = 2.8 m

3rd bright fringe, y3 = 4.7 cm = 0.047 m

a) let lamda is the wavelegnth of the laser light.

we know, y3 = 3*lamda*R/d

==> lamda = y3*d/(3*R)

= 0.047*0.12*10^-3/(3*2.8)

= 6.71*10^-7 m <<<<<<<<<<<<-------------------Answer

b) use, c = lamda*f

==> f = c/lamda

= 3*10^8/(6.71*10^-7)

= 4.47*10^14 Hz <<<<<<<<<<<<-------------------Answer


c) d*sin(theta) = 2*lamda

sin(theta) = 2*lamda/d

sin(theta) = 2*6.71*10^-7/(0.12*10^-3)

theta = sin^-1(0.01118)

= 0.641 degrees <<<<<<<<<<<<-------------------Answer